OpenStudy (anonymous):

In one popular game, five number cubes are rolled at the same time. If the number cubes are kept separate, what is the total possible number of outcome

OpenStudy (anonymous):

number cubes? you mean dice?

OpenStudy (anonymous):

six numbers on each "number cube" (good lord) so by the counting principle you have \[6\times 6\times 6\times 6\times 6=6^5\] possible outcomes